A common design for a swimming pool is for the depth to change gradually from the shallow end to the deep end. Use the dimensions shown to find the volume of water the pool can hold .
30ft
10ft-highest depth
4ft-lowest depth
15ft

Answer by checkley77(12844)

The dimensions of the pool are:
l*w*(average depth)
30*15(10+4)/2=450*7=3,150 ft^3 is the volume of the pool.